# Unexpected loss of maximality: the case of Hilbert square of real surfaces

### v3

@@ Abstract.
We explore the maximality of the Hilbert square of maximal real surfaces, and find that in many cases the Hilbert square is maximal if and only if the surface has connected real locus. In particular, the Hilbert square of no maximal $K3$-surface is maximal. Nevertheless, we exhibit maximal surfaces with disconnected real locus whose Hilbert square is maximal.
@@ 1. Introduction
A result of fundamental importance in understanding the topology of real algebraic varieties is the Smith inequality [26] that bounds from above the sum of the $\mathbb{F}_{2}$-Betti numbers of the fixed point set of an involution by the sum of $\mathbb{F}_{2}$-Betti numbers of the ambient space itself. Applied to the complex conjugation on a $n$-dimensional algebraic variety $X$ defined over the field $\mathbb{R}$ it states that
@@
In dimension one, examples of planar $M$-curves are given by the Harnack curves [15], and one can easily construct abstract maximal real algebraic curves of arbitrary genus. In higher dimensions, the question of sharpness is far from being completely understood despite the existence of a powerful patchworking method due to O. Viro. The latter one is presumed to provide maximal projective hypersurfaces of any dimension and degree, but yet only asymptotic (by degree) maximality is achieved [16]. Otherwise, apart from abelian varieties, flag manifolds equipped with the standard real structure, and smooth toric varieties [2], only sporadic examples are known.
@@
Let $X$ be a real nonsingular projective surface satisfying $H_{1}(X(\mathbb{C}),\mathbb{F}_{2})$ $=0$ and $h^{2,0}(X(\mathbb{C}))>0$, then $X^{[2]}$ is not maximal. In particular, this is the case if $X$ is a K3-surface, or a surface birational to it.
@@
If $X$ is a maximal real nonsingular rational surface with $X(\mathbb{R})$ connected, then $X^{[2]}$ is maximal. In particular, this is the case of the projective plane, and of ruled surfaces $\mathbb{P}_{\mathbb{P}^{1}}(\mathcal{E})$ where $\mathcal{E}$ is a rank 2 vector bundle equipped with a real structure that lifts the standard real structure on $\mathbb{P}^{1}.$
@@
Examples of maximal surfaces $X$ with torsion free homology, and such that $1<\beta_{0}(X(\mathbb{R}))<\beta_{1}(X)$ are given by the ruled surfaces $X=\mathbb{P}_{C}(E),$ where $E$ is a rank $2$ complex vector bundle over a maximal curve $C$ of positive genus. For such surfaces, the real locus $X(\mathbb{R})$ is disconnected, and we show that their Hilbert square is maximal, in contrast with situations described in Corollary 1.3. On the other hand, if $X$ is a maximal torus of complex dimension 2, we notice that $X^{[2]}$ is not maximal while $\beta_{0}(X(\mathbb{R}))=\beta_{1}(X(\mathbb{C})).$
@@
In all the enumerated results, as well as throughout the whole paper, when we speak about the maximality of a Hilbert scheme of points $X^{[n]},$ we mean the maximality of $X^{[n]}$ with respect to the real structure which is canonically inherited from a real structure on $X$. Thus, non maximality with respect to this kind of real structures does not exclude that $X^{[n]}$ may admit another real structure with respect to which it becomes maximal. In particular, at the current stage we do not know if there exist (real or complex) $K3,$ or Enriques, surfaces $X$ for which $X^{[n]}$ admits however a maximal real structure.
@@
All of the above definitions and the results, including their proofs, literally extend from real algebraic setting to compact complex analytic manifolds equipped either with an anti-holomorphic involution.
@@ 2.1. Smith theory
Most results cited in this section are due to P.A. Smith; proofs can be found, e.g., in [4, Chapter 3] and [6, Chapter 1].
Throughout the section we consider a topological space $X$ with a cellular involution $c:X\to X,$ i.e., $c$ transforms cells into cells and acts identically on each invariant cell. Denote by $F=\operatorname{Fix}c,\,\bar{X}=X/c,$ and let $\operatorname{in}:F\hookrightarrow X$ and $\operatorname{pr}:X\rightarrow\bar{X}$ be the natural inclusion and projection, respectively.
@@
[…]
Analyzing this sequence notice first that we have a canonical canonical splitting $\operatorname{Sm}_{*}(X)=S_{*}(F)\oplus\operatorname{Im}(1+c_{*}).$ The transfer homomorphism $\operatorname{tr}^{*}:S_{*}(\bar{X},F)\to\operatorname{Sm}_{*}(X,F)$ is an isomorphism [4, Chapter 3] (see also op. cit. for the cohomology version). In view of the above identifications the long exact sequences associated to (2.1) yields:
@@
The following immediate consequences of Theorem 2.1 which we state in the homology setting, have an obvious counterpart for cohomology:
@@ 2.2. Equivariant (co)homology and the Borel spectral sequence
The maximality of $(X,c)$ can be reformulated into the surjectivity of the restriction map from the equivariant cohomology to the usual cohomology (see, for example, [27, Chapter III, Proposition 4.16]):
@@ 3. Cut-and-Paste construction of Hilbert squares over the reals
For smooth varieties, a simple, well known construction of the Hilbert square is the following. Given a smooth variety $X$, one lifts the involution $\tau:X\times X\to X\times X$ permuting the factors to an involution $\operatorname{Bl}(\tau)$ on the blowup $\operatorname{Bl}_{\Delta}(X\times X)$ of $X\times X$ along the diagonal $\Delta\subset X\times X$. The quotient of $\operatorname{Bl}_{\Delta}(X\times X)$ by $\operatorname{Bl}(\tau)$ is then naturally isomorphic to the Hilbert square $X^{[2]}$.
The quotient map $\operatorname{Bl}_{\Delta}(X\times X)\to X^{[2]}$ is a ramified double covering. By construction, the branch locus $E\subset X^{[2]}$ is naturally isomorphic to $\mathbb{P}(T^{*}X)$ (in the Grothendieck notation) and coincides with the exceptional divisor of the canonical projection $X^{[2]}\to X^{(2)}$ to the symmetric square $X^{(2)}$ of $X.$ The normal bundle of $E$ in $X^{[2]}$ is naturally isomorphic to the square of the tautological bundle over $\mathbb{P}(T^{*}X)$.
This construction works over any field. Applying it to a smooth variety $X$ defined over the reals, we observe that $X^{[2]}(\mathbb{R})$ is a disjoint union of connected components
@@
$X^{[2]}(\mathbb{R})=X^{[2]}_{\text{main}}(\mathbb{R})\bigsqcup X^{[2]}_{\text{extra}}(\mathbb{R}),\quad X^{[2]}_{\text{extra}}(\mathbb{R})=\bigsqcup_{i<j}\left(F_{i}\times F_{j}\right)$
where $F_{1},\dots,F_{r}$ are the connected components of $X(\mathbb{R})$ and $X^{[2]}_{\text{main}}(\mathbb{R})$ is the component of $X^{[2]}(\mathbb{R})$ that contains $(X/\operatorname{c})\setminus X(\mathbb{R})$. If $X(\mathbb{R})=\emptyset$, then $X^{[2]}(\mathbb{R})=X^{[2]}_{\text{main}}(\mathbb{R})=X/\operatorname{c}$.
If $X(\mathbb{R})\neq\emptyset$ is not empty, then $E(\mathbb{R})\subset X^{[2]}_{\text{main}}(\mathbb{R})$ is also not empty and naturally diffeomorphic to $\mathbb{P}_{\mathbb{R}}(T^{*}X(\mathbb{R}))$. The normal line bundle of $E(\mathbb{R})$ in $X^{[2]}_{\text{main}}(\mathbb{R})$ is trivial, and $E(\mathbb{R})$ divides $X^{[2]}_{\text{main}}(\mathbb{R})$ in $r+1$ submanifolds with boundary:
$X^{[2]}_{\text{main}}(\mathbb{R})=\bigcup_{i=0}^{r}\mathbb{H}_{i},\quad\partial\mathbb{H}_{0}=\bigsqcup^{r}_{i=1}\partial\mathbb{H}_{i}$
where
@@
Here $\operatorname{c}:X\rightarrow X$ denotes the anti-holomorphic involution on $X$ induced the non-trivial element of $\operatorname{Gal}(\mathbb{C}/\mathbb{R}),$ while $\Delta F_{i}$ is the diagonal in $F_{i}^{(i)},$ for every $i=1,\dots,r.$ The manifolds $\mathbb{H}_{i}$ with $i=1,\dots,r$ are glued to $\mathbb{H}_{0}$ along common with $\mathbb{H}_{0}$ boundaries $\mathbb{P}_{\mathbb{R}}T^{*}F_{i}\subset\mathbb{P}_{\mathbb{R}}(T^{*}X(\mathbb{R}))$.
@@
    vanish at $\xi\in H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{k})$ if and only if $\xi\cap\omega=0$ where $\omega$ is the first Stiefel-Whitney class of the tautological line bundle over $\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$.
@@
    is non-vanishing on each $\xi\in H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{k})$ with $\xi\cap\omega\neq 0$; in particular, it is not vanishing on each of the fiber-classes $w_{k}\in H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{k})$.
@@
The given definition of $\omega$ is equivalent to saying that it is the characteristic class of the tautological double covering $UT^{*}F_{k}\to\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ (where $UT^{*}$ stands for the unit cotangent bundle). By the latter reason we may also interpret $\omega$ as (a restriction of) the characteristic class of the double coverings $X\setminus X(\mathbb{R})\to(X/\operatorname{c})\setminus X(\mathbb{R})$ and $(F_{k}\times F_{k})\setminus\Delta F_{k}\to(F_{k})^{(2)}\setminus\Delta F_{k},\,k=1,\dots,r.$ For that, it is sufficient, for each $k=1,\dots,r,$ to identify $UT^{*}F_{k}$ and $\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ with the boundary of a tubular neighborhood of $\Delta F_{k}$ in $F_{k}\times F_{k}$ and $(F_{k})^{(2)}$, respectively. Similarly, when it is a question of $X\setminus X(\mathbb{R})\to(X/\operatorname{c})\setminus X(\mathbb{R})$, we identify $\cup_{k}UT^{*}F_{k}$ and $\cup_{k}\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ with the boundary of a tubular neighborhood of $\cup F_{k}$ in $X$ and $X/\operatorname{c},$ respectively.
Due to $\omega$ being restricted from $H^{1}(\mathbb{H}_{k})$ for each $k=0,\dots,r$, the non-vanishing of $\mathfrak{i}_{*}^{\mathbb{H}_{k}}(\xi)$ when $\xi\cap\omega=1$ is evident. This proves the second statement of the lemma and the “only if part” of the first one.
To prove the “if part” of the first statement, let $k\geq 1$ and consider a loop $\Gamma\subset\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ representing a class $\xi$ with $\xi\cap\,\omega=0$. Such a loop lifts to a loop $\tilde{\Gamma}$ on the boundary of a tubular neighborhood of $\Delta F_{k}$ in $F_{k}\times F_{k}$. Therefore, the homology class $[\tilde{\Gamma}]\in H_{1}(F_{k}\times F_{k})$ belongs to the image of $H_{1}(\Delta F_{k})$. But, according to the Künneth formula and Poincaré duality, the class $[\tilde{\Gamma}],$ as every class in the image of $H_{1}(\Delta F_{k}),$ is of the form $\sum(x_{j}\otimes 1+1\otimes x_{j}),$ for some $x_{j}\in H_{1}(F_{k}).$ Since, in addition, the inclusion map
@@
$\mathfrak{i}_{*}^{\mathbb{H}_{k}}(\xi)=$ $\,(\operatorname{pr}_{*}\circ\operatorname{in}^{\complement\Delta F_{k}}_{*})(\tilde{\Gamma})$
@@
The first item follows, for example, from the more general formula computing the Euler characteristic of the Hilbert scheme of points of surfaces defined over the reals [20, page 5452], and the observation that $\chi(X)=\beta_{*}(X)-4\beta_{1}(X).$
The other two items are implicitly contained in the proof of Theorem 2.2 in [28]. According to [28, Theorem 2.2], $\operatorname{Tors}_{2}H_{*}(X^{[2]};\mathbb{Z})=0$ as soon as $\operatorname{Tors}_{2}H_{*}(X;\mathbb{Z})=0$. Therefore, both parts of the statement follow from the following relation for the ordinary Betti numbers
@@
[…]
@@
=\,= $\,\frac{1}{2}(2+2\beta_{1}+\beta_{2})^{2}+3+\beta_{1}+\frac{3}{2}\beta_{2}$
>\,>  $\,5+3\beta_{1}+\frac{1}{2}\beta_{2}$
[…]
@@
The map $\phi$ extends to the blowup $Bl_{p}(X)$ of $X$ at the point $p,$ and so we have a commutative diagram
${Bl_{p}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$$\scriptstyle{\operatorname{pr}}$${X^{[2]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\,}$$\scriptstyle{\pi}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${X^{(2)}.}$
@@
Note, first, that $H^{i}(Bl_{p}(X))=H^{i}(X)\oplus H^{i}(\mathbb{P}^{1})$ for any $i>0$. Here, the first summand coincide with $\operatorname{Im}\operatorname{pr}^{*}$. Since $f^{*}:H^{*}(X^{(2)})\rightarrow H^{*}(X)$ is surjective (see, for example, [11, Lemma 2.5] for a much more general statement), this implies that the first summand is contained in $\operatorname{Im}\phi^{*}$.
The second summand is nontrivial only for $i=2$. It is generated by $\mathbb{F}_{2}$-reduction of the integer class $\operatorname{in}^{*}c_{1}(\theta),$ where $\theta$ is the tautological line bundle over $\mathbb{P}(T^{*}X)=E\subset X^{[2]}$ and $\operatorname{in}:\mathbb{P}^{1}\hookrightarrow E$ is the inclusion. Since $c_{1}(\theta)$ is the restriction of an appropriate element $e\in H^{2}(X^{[2]};\mathbb{Z})$ (see, for example, [28, page 4]), the second summand is also contained $\operatorname{Im}\phi^{*}$. ∎

### v4

# On the maximality problem for the Hilbert square of real surfaces

@@ Abstract.
We explore maximality with respect to the classical Smith bound on the total Betti number of the real locus. For a large class of surfaces, we prove that the Hilbert square of a real surface is maximal if and only if the surface is maximal and has connected real locus. In particular, the Hilbert square of no K3 or abelian surface is maximal. We also exhibit various types of maximal surfaces, including ones with disconnected real locus, whose Hilbert square is maximal.
@@ 1. Introduction
A result of fundamental importance in understanding the topology of real algebraic varieties is the Smith inequality [29] that bounds from above the sum of the $\mathbb{F}_{2}$-Betti numbers of the fixed point set of an involution by the sum of $\mathbb{F}_{2}$-Betti numbers of the ambient space itself. Applied to the complex conjugation on an $n$-dimensional algebraic variety $X$ defined over the field $\mathbb{R}$ it states that
@@
In dimension one, examples of $M$-curves of arbitrary genus are given by Harnack [16]. In higher dimensions, the question of sharpness is far from being completely understood despite the existence of a powerful patchworking method due to O. Viro. The latter one is presumed to provide maximal projective hypersurfaces of any dimension and degree, but yet only asymptotic (by degree) maximality is achieved [17]. Otherwise, apart from abelian varieties, flag manifolds equipped with the standard real structure, and smooth toric varieties [2], only sporadic examples are known.
@@
Let $X$ be a real nonsingular projective surface satisfying $H_{1}(X(\mathbb{C}),\mathbb{F}_{2})$ =0=0 and $h^{2,0}(X(\mathbb{C}))>0$, then $X^{[2]}$ is not maximal. In particular, this is the case if $X$ is a K3-surface, or a surface birational to it.
@@
If $X$ is a maximal real nonsingular rational surface with $X(\mathbb{R})$ connected, then $X^{[2]}$ is maximal. In particular, this is the case for the projective plane and ruled surfaces $\mathbb{P}_{\mathbb{P}^{1}}(\mathcal{E}),$ where $\mathcal{E}$ is a rank 2 vector bundle equipped with a real structure that lifts the standard real structure on $\mathbb{P}^{1}.$
@@
Examples of maximal surfaces $X$ with torsion free homology and $1<\beta_{0}(X(\mathbb{R}))<\beta_{1}(X)$ are given by the ruled surfaces $X=\mathbb{P}_{C}(E),$ where $E$ is a rank $2$ complex vector bundle over a maximal curve $C$ of positive genus. For such surfaces, the real locus $X(\mathbb{R})$ is disconnected, and we show that their Hilbert square is maximal, in contrast with situations described in Corollary 1.3. On the other hand, if $X$ is a maximal torus of complex dimension 2, we notice that $X^{[2]}$ is not maximal while $\beta_{0}(X(\mathbb{R}))=\beta_{1}(X(\mathbb{C})).$
@@
In all of the enumerated results, as well as throughout the whole paper, when we speak about the maximality of a Hilbert scheme of points $X^{[n]},$ we mean the maximality of $X^{[n]}$ with respect to the real structure which is canonically inherited from a real structure on $X$. Thus, non maximality with respect to this kind of real structures does not exclude that $X^{[n]}$ may admit another real structure with respect to which it becomes maximal. In particular, at the current stage we do not know if there exist (real or complex) $K3,$ or Enriques, surfaces $X$ for which $X^{[n]}$ admits however a maximal real structure.
@@
All of the above definitions and the results, including their proofs, literally extend from real algebraic setting to compact complex analytic manifolds equipped with an anti-holomorphic involution.
@@ 2.1. Smith theory
Most of the results cited in this section are due to P.A. Smith; proofs can be found, e.g., in [4, Chapter 3] and [6, Chapter 1].
Throughout the section we consider a topological space $X$ equipped with a cellular involution $c:X\to X,$ i.e., an involution $c$ that transforms cells into cells and acts identically on each invariant cell. Denote by $F=\operatorname{Fix}c,\,\bar{X}=X/c,$ and let $\operatorname{in}:F\hookrightarrow X$ and $\operatorname{pr}:X\rightarrow\bar{X}$ be the natural inclusion and projection, respectively.
@@
$0\rightarrow\operatorname{Sm}_{*}(X)\xrightarrow{\text{inclusion}}S_{*}(X)\xrightarrow{1+c_{*}}\operatorname{Sm}_{*}(X,F)\rightarrow 0.$ (2.1)
Analyzing this sequence notice, first that we have a canonical canonical splitting $\operatorname{Sm}_{*}(X)=S_{*}(F)\oplus\operatorname{Im}(1+c_{*}).$ The transfer homomorphism $\operatorname{tr}^{*}:S_{*}(\bar{X},F)\to\operatorname{Sm}_{*}(X,F)$ is an isomorphism [4, Chapter 3] (see also op. cit. for the cohomology version). In view of the above identifications, the long exact sequences associated to (2.1) yield:
@@
The following immediate consequences of Theorem 2.1, which we state in the homology setting, have an obvious counterpart for cohomology:
@@ 2.2. Equivariant (co)homology and the Borel spectral sequence
The maximality of $(X,c)$ can be reformulated into the surjectivity of the restriction map from the equivariant cohomology to the usual cohomology (see, for example, [30, Chapter III, Proposition 4.16]):
@@ 3. Cut-and-Paste construction of Hilbert squares over the reals
Let $X$ be a nonsingular projective surface. By definition, the Hilbert square of $X,$ denoted by $X^{[2]},$ is the Hilbert scheme parametrizing 0-dimen-sional subschemes of $X$ of length $2.$ As it follows from the universal property of the Hilbert schemes, $X^{[2]}$ has an elementary description (see, e.g., [9, Example 7.3.1]), which we recall for the convenience of the reader.
@@
Let $\operatorname{Bl}_{\Delta}(X\times X)$ be the blow-up of $X\times X$ along the diagonal $\Delta\subset X\times X.$ The involution $\tau$ on $X\times X$ permuting the factors lifts to an involution $\operatorname{Bl}(\tau)$ on $\operatorname{Bl}_{\Delta}(X\times X).$ The quotient of $\operatorname{Bl}_{\Delta}(X\times X)$ by $\operatorname{Bl}(\tau)$ is naturally isomorphic to the Hilbert square $X^{[2]}.$ The fixed locus of $\operatorname{Bl}(\tau)$ is the exceptional divisor of the blow-up $\operatorname{Bl}_{\Delta}(X\times X).$ The smoothness of this divisor implies that $X^{[2]}$ is nonsingular.
The branch locus $E\subset X^{[2]}$ of the double ramified covering $\operatorname{Bl}_{\Delta}(X\times X)\to X^{[2]}$ is naturally isomorphic to the exceptional divisor of the blowup, and, since the normal vector bundle of $\Delta$ in $X\times X$ is isomorphic to the tangent vector bundle $TX$ of $X,$ both this exceptional divisor and $E$ are naturally isomorphic to $\mathbb{P}(T^{*}X)$ [^1] In agreement with [31], which we use as reference, we follow the Grothendieck convention according to which the projectivization of a vector space is the space of its hyperplanes.. In other words, a point of $E$ is identified with a point of $X$ plus a complex line in the tangent space at that point. Pairs of distinct points in $X$ represent the points of the complement, $X^{[2]}\setminus E.$
Notice that the normal bundle of $E$, as a branch locus, is the square of the normal bundle of the ramification locus. Hence, it is naturally isomorphic to the square of the tautological bundle over $\mathbb{P}(T^{*}X)$.
The blowing-up map $\operatorname{Bl}_{\Delta}(X\times X)\rightarrow X\times X$ descends to a morphism,
$\pi:X^{[2]}\rightarrow X^{(2)},$ (3.1)
where $X^{(2)}=(X\times X)/\tau$ is the $2$-fold symmetric product. This morphism is called the Hilbert-Chow map. Notice that the map $\pi$ is one to one on $X^{[2]}\setminus E$ and maps $E$ to the diagonal $\Delta X$ of $X^{(2)}.$
The construction of the Hilbert square is independent on the choice of the ground field. Applying it to a real nonsingular projective surface $(X,\operatorname{c}),$ it equips $X^{[2]}$ with a natural real structure, still denoted by $\operatorname{c}$. It acts on points in $X^{[2]}$ represented by pairs of distinct points in $X$ by sending a pair to the complex conjugate one, while if a point of $X^{[2]}$ is represented by a point of $X$ with a tangent line, then it is sent to the conjugate point with the conjugate line.
Thus, $X^{[2]}(\mathbb{R})=X/\operatorname{c}$ if $X(\mathbb{R})=\emptyset,$ and otherwise $X^{[2]}(\mathbb{R})$ is a smooth manifold with $\mbox{dim}_{\mathbb{R}}X^{[2]}(\mathbb{R})=4,$ having the following properties:
  • $E(\mathbb{R})\subset X^{[2]}(\mathbb{R})$ is a submanifold of real codimension 1 in $X(\mathbb{R}).$ It is diffeomorphic to
    $\mathbb{P}_{\mathbb{R}}(T^{*}X(\mathbb{R}))=\bigsqcup_{i=1}^{r}\mathbb{P}_{\mathbb{R}}(T^{*}F_{i}),$
    where $F_{1},\dots,F_{r}$ are the connected components of $X(\mathbb{R}).$
  • $X^{[2]}(\mathbb{R})\setminus E(\mathbb{R})$ is the disjoint union of the connected real $4$-dimensional manifolds
    $\left((X/\operatorname{c})\setminus X(\mathbb{R})\right)\sqcup\bigsqcup_{i=1}^{r}\left(F_{i}^{(2)}\setminus\Delta F_{i}\right)\sqcup\bigsqcup_{1\leq i<j\leq r}F_{i}\times F_{j},$
    where $\Delta F_{i}$ stands for the diagonal in $F_{i}^{(2)}.$ Each of the manifolds $F_{i}\times F_{j}$ in this decomposition is closed, while the manifolds $(X/\operatorname{c})\setminus X(\mathbb{R})$ and $F_{i}^{(2)}\setminus\Delta F_{i},\,i=1,\dots,r$ are open.
  • For each $1\leq i\leq r$, the component $\mathbb{P}_{\mathbb{R}}(T^{*}F_{i})$ of $E(\mathbb{R})$ is a common boundary of $(X/\operatorname{c})\setminus X(\mathbb{R})$ with $F_{i}^{(2)}\setminus\Delta F_{i}$.
From this we can conclude that $X^{[2]}(\mathbb{R})$ has the following decomposition, as illustrated in the figure below:
$\sqcup$$\bigsqcup_{1\leq i<j\leq r}$$\mathbb{H}_{0}$$\mathbb{H}_{1}$$\mathbb{H}_{2}$$\mathbb{H}_{r}$$\mathbb{P}_{\mathbb{R}}(T^{*}F_{r})$$\mathbb{P}_{\mathbb{R}}(T^{*}F_{2})$$\mathbb{P}_{\mathbb{R}}(T^{*}F_{1})$$F_{i}\times F_{j}$$X^{[2]}(\mathbb{R})_{\rm extra}$$X^{[2]}(\mathbb{R})_{\rm main}$
Figure 1. The decomposition of $X^{[2]}(\mathbb{R})$
@@
$X^{[2]}(\mathbb{R})=X^{[2]}_{\text{main}}(\mathbb{R})\bigsqcup X^{[2]}_{\text{extra}}(\mathbb{R}),\quad X^{[2]}_{\text{extra}}(\mathbb{R})=\bigsqcup_{1\leq i<j\leq r}\left(F_{i}\times F_{j}\right),$
[…]
$X^{[2]}_{\text{main}}(\mathbb{R})=\bigcup_{i=0}^{r}\mathbb{H}_{i},$
such that
@@
Each manifold $\mathbb{H}_{i},\,i=1,\dots,r$ is glued to $\mathbb{H}_{0}$ along their common boundary $\mathbb{P}_{\mathbb{R}}(T^{*}F_{i})\subseteq\mathbb{P}_{\mathbb{R}}(T^{*}X(\mathbb{R})).$
Notice that the Hilbert-Chow map (3.1) is equivariant with respect to the induced real structures on $X^{[2]}$ and $X^{(2)}$.
@@
    vanish at $\xi\in H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{k})$ if and only if $\xi\cap\omega=0,$ where $\omega$ is the first Stiefel-Whitney class of the tautological line bundle over $\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$.
@@
    is non-vanishing on each $\xi\in H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{k})$ with $\xi\cap\omega\neq 0.$ In particular, it is not vanishing on each of the fiber-classes $w_{k}\in H_{1}(\mathbb{P}_{\mathbb{R}}T^{*}F_{k}).$
@@
The given definition of $\omega$ is equivalent to saying that it is the characteristic class of the tautological double covering $UT^{*}F_{k}\to\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ (where $UT^{*}$ stands for the unit cotangent bundle). By the latter reason, we may also interpret $\omega$ as a restriction of the characteristic class $v_{k}$ of the double covering $(F_{k}\times F_{k})\setminus\Delta F_{k}\to(F_{k})^{(2)}\setminus\Delta F_{k},$ or as a restriction of the characteristic class $v$ of the double covering $X\setminus X(\mathbb{R})\to(X/\operatorname{c})\setminus X(\mathbb{R})$. For the former one, it is sufficient to identify $UT^{*}F_{k}$ and $\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ with the boundary of a tubular neighborhood of $\Delta F_{k}$ in $F_{k}\times F_{k}$ and $(F_{k})^{(2)}$, respectively. Similarly, when it is a question of $X\setminus X(\mathbb{R})\to(X/\operatorname{c})\setminus X(\mathbb{R})$, we identify $\cup_{k}UT^{*}F_{k}$ and $\cup_{k}\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ with the boundary of a tubular neighborhood of $\cup F_{k}$ in $X$ and $X/\operatorname{c},$ respectively. For more details on extending the class $\omega$ to cohomology classes in $\mathbb{H}_{0}$ and $\mathbb{H}_{i},\,i=1,\dots,r,$ we refer the interested reader to [22, Section 3.1].
Therefore, for each $\xi\in H^{1}(\mathbb{H}_{k})$ with $\xi\cap\omega=1,$ by the projection formula, we have
$\mathfrak{i}_{*}^{\mathbb{H}_{k}}(\xi)\cap v_{k}=\xi\cap(\mathfrak{i}^{\mathbb{H}_{k}})^{*}(v_{k})=\xi\cap\omega=1.$
This implies the “only if part” of the first statement and the second statement.
To prove the “if part” of the first statement, let $k\geq 1$ and consider a loop $\Gamma\subset\mathbb{P}_{\mathbb{R}}T^{*}F_{k}$ representing a class $\xi$ with $\xi\cap\,\omega=0$. Such a loop lifts to a loop $\tilde{\Gamma}$ on the boundary of a tubular neighborhood of $\Delta F_{k}$ in $F_{k}\times F_{k}$. Therefore, the homology class $\tilde{\Gamma}]\in H_{1}(F_{k}\times F_{k})$ belongs to the image of $H_{1}(\Delta F_{k})$. But, according to the Künneth formula and Poincaré duality, the class $\tilde{\Gamma}$], as every class in the image of $H_{1}(\Delta F_{k}),$ is of the form $\sum(x_{j}\otimes 1+1\otimes x_{j}),$ for some $x_{j}\in H_{1}(F_{k}).$ Since, in addition, the inclusion map
@@
[…]
@@
Corollary 3.2.
For each $k=1,\dots,r,$ we have $\ker\mathfrak{i}_{*}^{\mathbb{H}_{0}}\subseteq\ker\mathfrak{i}_{*}^{\mathbb{H}_{k}}.$
Proof.
The inclusion $\ker\mathfrak{i}_{*}^{\mathbb{H}_{0}}\subseteq\ker\mathfrak{i}_{*}^{\mathbb{H}_{k}}$ follows from the first part of the second statement and the ”if part” of the first statement of Proposition 3.1. ∎
@@
The first item follows, for example, from the more general formula computing the Euler characteristic of the Hilbert scheme of points of surfaces defined over the reals [21, page 5452], and the observation that $\chi(X)=\beta_{*}(X)-4\beta_{1}(X).$
The other two items are implicitly contained in the proof of Theorem 2.2 in [31]. According to [31, Theorem 2.2], $\operatorname{Tors}_{2}H_{*}(X^{[2]};\mathbb{Z})=0$ as soon as $\operatorname{Tors}_{2}H_{*}(X;\mathbb{Z})=0$. Therefore, both parts of the statement follow from the following relation for the ordinary Betti numbers
@@
This relation follows from the presentation of $X^{[2]}$ as the quotient of the blowup $\operatorname{Bl}_{\Delta}(X\times X)$ of X×X $\times X$ due to the following arguments:
@@
$=\,$ $\,\frac{1}{2}(2+2\beta_{1}+\beta_{2})^{2}+3+\beta_{1}+\frac{3}{2}\beta_{2}$
$>\,$  $\,5+3\beta_{1}+\frac{1}{2}\beta_{2}$
$\geq\,$ $\,\beta_{*}(X^{[2]}(\mathbb{R})),$
@@
The map $\phi$ extends to the blowup $\operatorname{Bl}_{p}(X)$ of $X$ at the point $p,$ and so we have a commutative diagram
${\operatorname{Bl}_{p}(X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi}$$\scriptstyle{\operatorname{pr}}$${X^{[2]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\,}$$\scriptstyle{\pi}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${X^{(2)}.}$
@@
Note, first, that $H^{i}(\operatorname{Bl}_{p}(X))=H^{i}(X)\oplus H^{i}(\mathbb{P}^{1})$ for any $i>0$. Here, the first summand coincide with $\operatorname{Im}\operatorname{pr}^{*}$. Since $f^{*}:H^{*}(X^{(2)})\rightarrow H^{*}(X)$ is surjective (see, for example, [12, Lemma 2.5] for a more general statement), this implies that the first summand is contained in $\operatorname{Im}\phi^{*}$.
The second summand is nontrivial only for $i=2$. It is generated by $\mathbb{F}_{2}$-reduction of the integer class $\operatorname{in}^{*}c_{1}(\theta),$ where $\theta$ is the tautological line bundle over $\mathbb{P}(T^{*}X)=E\subset X^{[2]}$ and $\operatorname{in}:\mathbb{P}^{1}\hookrightarrow E$ is the inclusion. Since $c_{1}(\theta)$ is the restriction of an appropriate element $e\in H^{2}(X^{[2]};\mathbb{Z})$ (see, for example, [31, page 4]), the second summand is also contained $\operatorname{Im}\phi^{*}$. ∎